# Analysis of Shuffling Beyond Pure Local Differential Privacy

### v2

@@ B.2 $k$-$R$
Lower shuffle index. For any distinct pair $x_{1}\neq x_{1}^{\prime}\in[k]$, the privacy amplification random variable with respect to the blanket distribution is
@@
so that $\mathbb{E}[l_{0}(Y)]=0$ and $\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime}):=\mathrm{Var}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})\bigr)=2k(p-q)^{2}.$ By definition,
$\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})=\sqrt{\frac{\gamma}{\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime})}}=\sqrt{\frac{q}{2(p-q)^{2}}}=\sqrt{\frac{\mathrm{e}^{\varepsilon_{0}}+k-1}{2\bigl(\mathrm{e}^{\varepsilon_{0}}-1\bigr)^{2}}}.$
In particular, $\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$ does not depend on the particular choice of the distinct pair $(x_{1},x_{1}^{\prime})$, and hence
$\underline{\chi}_{\mathrm{lo}}=\inf_{x_{1}\neq x_{1}^{\prime}}\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})=\sqrt{\frac{q}{2(p-q)^{2}}}.$
Upper shuffle index. For the upper shuffle index we take the reference distribution to be one of the local randomizers $\mathcal{R}_{x}$. Fix $x_{1}\neq x_{1}^{\prime}$, and consider
@@
$\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})=\inf_{x\in[k]}\sqrt{\frac{1}{\mathrm{Var}(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x}))}}=\sqrt{\frac{q}{2(p-q)^{2}}}$
Thus, $\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})$ also does not depend on the particular choice of the distinct pair $(x_{1},x_{1}^{\prime})$, and hence, for $k$-randomized response with $k\geq 3$, the lower and upper shuffle indices coincide:
$\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}=\sqrt{\frac{q}{2(p-q)^{2}}}=\sqrt{\frac{\mathrm{e}^{\varepsilon_{0}}+k-1}{2\bigl(\mathrm{e}^{\varepsilon_{0}}-1\bigr)^{2}}}.$
Appendix C Implementations of FFT algorithms
C.1 Parameters

### v3

@@ C.2 $k$-$R$
C.2.1 Lower shuffle index.
For any distinct pair $x_{1}\neq x_{1}^{\prime}\in[k]$, the privacy amplification random variable with respect to the blanket distribution is
@@
so that $\mathbb{E}[l_{0}(Y)]=0$ and $\sigma_{\mathrm{BG}}^{2}:=\mathrm{Var}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})\bigr)=2k(p-q)^{2}.$ By definition,
$\chi=\sqrt{\frac{\gamma}{\sigma_{\mathrm{BG}}^{2}}}=\sqrt{\frac{q}{2(p-q)^{2}}}=\sqrt{\frac{\mathrm{e}^{\varepsilon_{0}}+k-1}{2\bigl(\mathrm{e}^{\varepsilon_{0}}-1\bigr)^{2}}}.$
In particular, $\chi$ does not depend on the particular choice of the distinct pair $(x_{1},x_{1}^{\prime})$, and hence
$\chi_{\mathrm{lo}}=\inf_{x_{1}\neq x_{1}^{\prime}}\chi=\sqrt{\frac{q}{2(p-q)^{2}}}.$
C.2.2 Upper shuffle index.
For the upper shuffle index we take the reference distribution to be one of the local randomizers $\mathcal{R}_{x}$. Fix $x_{1}\neq x_{1}^{\prime}$, and consider
@@
$\chi_{\mathrm{up}}=\inf_{x\in[k]}\sqrt{\frac{1}{\mathrm{Var}(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x}))}}=\sqrt{\frac{q}{2(p-q)^{2}}}$
Thus, $\chi_{\mathrm{up}}$ also does not depend on the particular choice of the distinct pair $(x_{1},x_{1}^{\prime})$, and hence, for $k$-randomized response with $k\geq 3$, the lower and upper shuffle indices coincide:
$\chi_{\mathrm{up}}=\chi_{\mathrm{lo}}=\sqrt{\frac{q}{2(p-q)^{2}}}=\sqrt{\frac{\mathrm{e}^{\varepsilon_{0}}+k-1}{2\bigl(\mathrm{e}^{\varepsilon_{0}}-1\bigr)^{2}}}.$
C.2.3 Comparison to the Clone Paradigm
Here, we consider Theorem 7.4 of Feldman [15]: the specific clone reduction of kkRR. Let the categorical output alphabet be $\mathcal{A}:=\{0,1,2,3\}$. Fix parameters $p=\frac{1}{e^{\varepsilon_{0}}+k-1}$ and $q=\frac{k-2}{e^{\varepsilon_{0}}+k-1}$. Define the reference categorical distribution $Q_{\mathrm{ref}}$ on $\mathcal{A}$ by
$Q_{\mathrm{ref}}(0)=p,\qquad Q_{\mathrm{ref}}(1)=p,\qquad Q_{\mathrm{ref}}(2)=q,\qquad Q_{\mathrm{ref}}(3)=1-2p-q.$
Define the two target categorical distributions $Q_{0},Q_{1}$ on $\mathcal{A}$:
$Q_{0}(0)=e^{\varepsilon_{0}}p,\qquad Q_{0}(1)=p,\qquad Q_{0}(2)=1-(1+e^{\varepsilon_{0}})p,\qquad Q_{0}(3)=0,$
$Q_{1}(0)=p,\qquad Q_{1}(1)=e^{\varepsilon_{0}}p,\qquad Q_{1}(2)=1-(1+e^{\varepsilon_{0}})p,\qquad Q_{1}(3)=0.$
The reduced shuffled mechanism is then
$\widehat{\mathcal{M}}_{b}\;:=\;\mathcal{S}\bigl(Z_{1}^{(b)},Z_{2},\dots,Z_{n}\bigr),\qquad Z_{1}^{(b)}\sim Q_{b},\ \ Z_{2},\dots,Z_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}Q_{\mathrm{ref}},\qquad b\in\{0,1\}.$
That is, one target user draws a single categorical message from a distribution $Q_{b}$, while each of the remaining n1n-1 users draws i.i.d. messages from a reference categorical distribution $Q_{\mathrm{ref}}$; all $n$ messages are then passed through an ideal shuffler. From the discussion in Section 3.1, the hockey-stick divergence between the two shuffled outputs can be expressed as the blanket divergence with $\gamma=1$. That is, we can interpret that Theorem 7.4 states that the blanket divergence upper boundeds the hockey-stick divergence of shuffled kkRR. Under this interpretation, we may analyze the reduced categorical instance with $\gamma=1$ and identify the corresponding shuffle index.
Writing $a:=e^{\varepsilon_{0}}$, we obtain the pointwise values of the privacy amplification random variable $l_{0}(y)$:
$l_{0}(0)=\frac{ap-p}{p}=a-1,\qquad l_{0}(1)=\frac{p-ap}{p}=1-a=-(a-1),\qquad l_{0}(2)=0,\qquad l_{0}(3)=0,$
since $Q_{0}(2)=Q_{1}(2)$ and $Q_{0}(3)=Q_{1}(3)=0$. Therefore, under $Y\sim Q_{\mathrm{ref}}$,
$l_{0}(Y)=\begin{cases}+(a-1)&\text{w.p. }p,\\ -(a-1)&\text{w.p. }p,\\ 0&\text{w.p. }1-2p,\end{cases}$
which implies
$\sigma^{2}=\mathrm{Var}(l_{0}(Y))=\mathbb{E}[l_{0}(Y)^{2}]=2p\,(a-1)^{2}.$
Consequently, the shuffle index for this $\gamma=1$ categorical instance is
$\chi\;=\;\frac{1}{\sqrt{2p}\,(a-1)}.$
Finally, one has $p=\frac{1}{a+k-1}$ (equivalently, $p=\frac{1}{e^{\varepsilon_{0}}+k-1}$), hence
$\chi\;=\;\frac{\sqrt{e^{\varepsilon_{0}}+k-1}}{\sqrt{2}\,\bigl(e^{\varepsilon_{0}}-1\bigr)}.$
This corresponds to $\chi_{\mathrm{lo}}$ of kkRR computed in the previous section.
Appendix D Implementations of FFT algorithms
D.1 Parameters